# Accurate estimation of the normalized mutual information of multidimensional data

### v1

@@ II.1 Normalizing the mutual information
rendering the differential and relative definitions equivalent. It is important to stress, however, that this equivalence does not hold for the entropies, i.e., $H_{\rm d}(X)\neq H_{\rm r}(X)$, and therefore also not for the NMI in Eq. (8). Hence, it is crucial to apply relative entropies [Eq. (11)] when we compute the normalization between two continuous random variables. Otherwise, the usage of different representations or parametrizations of the two probability densities could lead to different results for the NMI, which renders it challenging to compare or interpret the results.
@@
where $d_{X}$ and $d_{Y}$ denote the dimension of variables $X$ and $Y$, respectively. The definition of $I_{\text{GY}}(X,Y)$ clearly provides a mapping from the $I(X,Y)\in[0,\infty)$ to a normalized quantity. Remarkably, in the special case that $P(x,y)$ is a bivariate normal distribution, $I_{\text{GY}}$ can be shown to be identical to the linear correlation $|\rho|$. In the general case, however, the interpretation of $I_{\text{GY}}$ as a correlation measure is less clear. In particular, $I_{\text{GY}}$ shows relatively large values already for weakly correlated data.Diez22
@@ II.2 KSG-estimator using the continuous entropy
While various continuous entropy estimators have been proposed,Nemenman01 ; Jiao15 ; Archer14 ; Valiant17 we wish to adapt the KSG-estimatorKraskov04 to compute the NMI, as it is able to treat multidimensional problems. The algorithm is based on local $k$-nn statistics, that is, it aims to calculate the probability distribution $p_{i}$ of the distance between the data points $\{x_{i}\}$ and their $k$th nearest neighbor, assuming a local constant density. The entropy is then given as the expectation value of $\ln p_{i}$, which can be calculated from the digamma function $\psi$. For further reference, we only mention here the main results of the algorithm, and refer to the original paper for details.Kraskov04
Starting with the definition of differential entropy in Eq. (9), Kraskov et al.Kraskov04 derived unbiased estimators $\hat{H}_{\rm d}$ for the differential entropy and the joint entropy
@@
$+(d_{X}+d_{Y})\langle\ln 2\varepsilon\rangle,$ (19)
@@ II.3 Estimating the Invariant Measures
where $\tilde{\varepsilon}=\varepsilon/\langle\varepsilon^{d_{X}+d_{Y}}\rangle^{1/(d_{X}+d_{Y})}$ denotes the scaling-invariant $k$-nn radius. Equations (29)–(32) represent the main theoretical results of this paper. As the inequalities in Eq. (7) apply for the above entropy estimators $\hat{H}_{r}$, the NMI can be obtained through the normalization factor $(\hat{H}_{r}(X)\hat{H}_{r}(Y))^{-\frac{1}{2}}$ as in Eq. (8), or alternatively via the factor $\hat{H}_{r}(X,Y)^{-1}$ as in Eq. (16).
@@ III.2 Comparison of correlation measures
That is, due to the scalar product in Eq. (1), only diagonal elements of $\rho_{ij}^{\alpha\beta}$ with $\alpha\!=\!\beta$ occur. Comparing the results for the Pearson coefficient and the NMI, Figs. 3c and 3d show that the two correlation measures differ substantially, i.e., the structure of the Pearson matrix shows prominent patterns that are absent in the NMI. This artifact of the Pearson coefficient is due to the fact that the components $\rho_{ij}^{\alpha\alpha}$ at the same time account for the correlation between residues $i$ and $j$ and the relative orientation of the directions $\alpha=x,y,z$. Since the components can take positive and negative values ($-1\leq\rho_{ij}^{\alpha\alpha}\leq 1$), for example, a large positive value for $\rho_{ij}^{xx}$ can be canceled by a large negative value for $\rho_{ij}^{yy}$, although residues $i$ and $j$ are highly correlated. Showing the matrices of the three components $\rho_{ij}^{xx}$, $\rho_{ij}^{yy}$, and $\rho_{ij}^{zz}$, Fig. S4 readily explains the spurious pattern of the Pearson correlation matrix. Although this failure of the Pearson measure for Cartesian coordinates was noticed previously, Ichiye91 ; Lange06 the Pearson correlation coefficient nevertheless has seen widespread use in the construction of allosteric networks.Sethi09 ; McClendon09 ; Bhattacharyya16 ; Wodak19
@@
This leads to an eigenvalue problem for the canonical variables $\widehat{X}$ and $\widehat{Y}$, which yields the canonical correlation from a sum of the eigenvalues, i.e., $\rho_{\rm C}^{2}=\frac{1}{3}\sum_{i}\lambda_{i}$. As shown in the SI, the eigenvalues can be calculated directly from the $3\times 3$ correlation matrices, which renders CCA a rather efficient method. Comparing the resulting CCA matrix to the NMI, Fig. S5b reveals good overall agreement of the two correlation measures. This shows that the main discrepancy between the Pearson correlation and the NMI is due to the lack of an appropriate treatment of the relative orientation of the components, rather than intrinsically nonlinear correlation effects. Lange06

### v2

@@ II.1 Normalizing the mutual information
rendering the differential (subscript ’d’) and relative (subscript ’r’) definitions equivalent. It is important to stress, however, that this equivalence does not hold for the entropies, i.e., $H_{\rm d}(X)\neq H_{\rm r}(X)$, and therefore also not for the NMI in Eq. (8). Hence, it is crucial to apply relative entropies [Eq. (11)] when we compute the normalization between two continuous random variables. Otherwise, the usage of different representations or parametrizations of the two probability densities could lead to different results for the NMI, which renders it challenging to compare or interpret the results.
@@
where $d_{X}$ and $d_{Y}$ denote the dimension of variables $X$ and $Y$, respectively. The definition of $I_{\text{GY}}(X,Y)$ clearly provides a mapping from the $I(X,Y)\in[0,\infty)$ to a normalized quantity. While this can be achieved by various mappings, $I_{\text{GY}}$ can be shown to be identical to the linear correlation $|\rho|$ in the special case that $P(x,y)$ is a bivariate normal distribution. In the general case, however, the interpretation of $I_{\text{GY}}$ as a correlation measure is less clear, because it does not take into account the upper bound of the mutual information based on the individual entropies of $H(X)$ and $H(Y)$, see Eq. (7). In particular, $I_{\text{GY}}$ shows relatively large values already for weakly correlated data.Diez22
@@ II.2 KSG-estimator using the continuous entropy
While various continuous entropy estimators have been proposed,Nemenman01 ; Jiao15 ; Archer14 ; Valiant17 we wish to adapt the KSG-estimatorKraskov04 to compute the NMI, as it is able to treat multidimensional problems. The algorithm is based on local $k$-nn statistics, that is, it aims to calculate the probability distribution $p_{i}$ of the distance between the data points $\{x_{i}\}$ and their $k$th nearest neighbor, assuming a local constant density. This probability can be visualized as the volume of a $d$-dimensional sphere with radius $2\varepsilon_{i}$, which is a proxy for the density of the points around $x_{i}$. The entropy is then given as the expectation value of $-\ln p_{i}$, which can be calculated from the digamma function $\psi$. For further reference, we only mention here the main results of the algorithm, and refer to the original paper for details.Kraskov04
Kraskov et al.Kraskov04 used the multinomial theorem to calculate the probability $p_{i}$ and averaged the log-likelihood over all particles to obtain an estimate of the entropy. Starting with the definition of differential entropy in Eq. (9), they derived unbiased estimators $\hat{H}_{\rm d}$ for the differential entropy and the joint entropy
@@
$+(d_{X}+d_{Y})\langle\ln 2\varepsilon\rangle.$ (19)
@@ II.3 Estimating the Invariant Measures
where $\tilde{\varepsilon}=\varepsilon/\langle\varepsilon^{d_{X}+d_{Y}}\rangle^{1/(d_{X}+d_{Y})}$ denotes the scaling-invariant $k$-nn radius. Equations (29)–(32) represent the main theoretical results of this paper. As the inequalities in Eq. (7) apply for the above entropy estimators $\hat{H}_{r}$, the NMI can be obtained through the normalization factor $(\hat{H}_{r}(X)\hat{H}_{r}(Y))^{-\frac{1}{2}}$ as in Eq. (8), or alternatively via the factor $\hat{H}_{r}(X,Y)^{-1}$ as in Eq. (16). We note that Eqs. (29)–(32) address the two main challenges in computing MI: accurately estimating probability distributions in high-dimensional spaces with finite data, and normalizing MI for meaningful comparisons across different data sets. The first challenge is addressed by using $k$-nearest neighbor statistics to approximate probability distributions; the second one by normalizing MI using relative entropy, which is achieved by introducing an invariant measure that assumes a constant density.
@@ III.2 Comparison of correlation measures
That is, due to the scalar product in Eq. (1), only diagonal elements of $\rho_{ij}^{\alpha\beta}$ with $\alpha\!=\!\beta$ occur. Comparing the results for the Pearson coefficient and the NMI, Figs. 3c and 3d show that the two correlation measures differ substantially, i.e., the structure of the Pearson matrix shows prominent patterns that are absent in the NMI. This artifact of the Pearson coefficient is due to the fact that the components $\rho_{ij}^{\alpha\alpha}$ at the same time account for the correlation between residues $i$ and $j$ and the relative orientation of the directions $\alpha=x,y,z$. Since the components can take positive and negative values ($-1\leq\rho_{ij}^{\alpha\alpha}\leq 1$), for example, a large positive value for $\rho_{ij}^{xx}$ can be canceled by a large negative value for $\rho_{ij}^{yy}$, although residues $i$ and $j$ are highly correlated. Showing the matrices of the three components $\rho_{ij}^{xx}$, $\rho_{ij}^{yy}$, and $\rho_{ij}^{zz}$, Fig. S4 readily explains the spurious pattern of the Pearson correlation matrix. Although this failure of the Pearson measure for Cartesian coordinates was previously pointed out by Lange and Grubmüller, Lange06 the Pearson correlation coefficient nevertheless has seen widespread use in the construction of allosteric networks.Sethi09 ; McClendon09 ; Bhattacharyya16 ; Wodak19
@@
This leads to an eigenvalue problem for the canonical variables $\widehat{X}$ and $\widehat{Y}$, which yields the canonical correlation from a sum of the eigenvalues, i.e., $\rho_{\rm C}^{2}=\frac{1}{3}\sum_{i}\lambda_{i}$. As shown in the SI, the eigenvalues can be calculated directly from the $3\times 3$ correlation matrices, which renders CCA a rather efficient method. Comparing the resulting CCA matrix to the NMI, Fig. S5b reveals good overall agreement of the two correlation measures. As first discussed by Lange and Grubmüller,Lange06 this shows that the main discrepancy between the Pearson correlation and the NMI is due to the lack of an appropriate treatment of the relative orientation of the components, rather than intrinsically nonlinear correlation effects.